Calculate the volume in milliliters of a 0.51M barium chloride solution that contains 75.0mmol of barium chloride BaCl2. Be sure your answer has the correct number of significant digits.
Answer #1

Formula: molarity of a solution = no. of mmol of the solute/volume (mL) of the solution

i.e. 0.51 mmol/mL = 75 mmol/volume of solution

Therefore, the volume of barium chloride solution = (75/0.51) mL = 147.1 mL = 1.471*102~ 1.5*102 mL

Note: The answer has two significant digits.